OpenStudy (anonymous):

would 2/3x; 4 units down = y=2/3x-4?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

it would be y= 2/3(x-4)?

OpenStudy (anonymous):

*wouldnt

OpenStudy (mertsj):

it would be y= 2/3(x-4)? NO!!!!!

OpenStudy (anonymous):

huh?

OpenStudy (mertsj):

y=2/3x-4? This one is correct. If you put the parentheses around it, you have moved it 4 to the right.

OpenStudy (anonymous):

Oh ok got you.

OpenStudy (mertsj):

Good

OpenStudy (anonymous):

the correct answer is \[y=\frac{2}{3x}\] 4 unite down is \[y=\frac{2}{3x}-4\]